OpenStudy (anonymous):

Solve for x in the following equation -2x^3 = 128

OpenStudy (anonymous):

@navk

OpenStudy (anonymous):

First step will be to divide by -2 on both sides of the equation. That will give you x^3 = -64 Now take the cube root..

OpenStudy (anonymous):

divide the cube root by -2?

OpenStudy (anonymous):

no divide 128 by -2 to get -64 then find the cube root of -64

OpenStudy (solomonzelman):

\[-64=(-4)^3\]

OpenStudy (solomonzelman):

Overview: 1) divide both sides by -2 2) cube root both sides

OpenStudy (anonymous):

so it would be x = -4i since you cube root both sides?

OpenStudy (anonymous):

Cube root of -64 is not -4i the imaginary number 'i' appears mostly in square roots not cube roots

OpenStudy (anonymous):

Thats because -4 times -4 times -4 = -64 on the other hand -4i times -4i times -4i = -64i^3 = -64 * i^2 * i = -64 * -1 * i = 64i So -4i can not be a cube root of -64

OpenStudy (anonymous):

x = -4

OpenStudy (solomonzelman):

Yes, there are never imaginary numbers when \(\LARGE\color{blue}{ \bf \sqrt[odd]{something} }\)

OpenStudy (solomonzelman):

No imaginaries. you know that \(\large\color{blue}{ \bf -64=(-4)^3 }\) AND \(\large\color{blue}{ \bf x^3=x }\)